OpenStudy (anonymous):

Help, please. Partial Derivatives when it comes with more than two variables. Let say a function f(x,y,z)

hartnn (hartnn):

\(\partial f/\partial x\) means you are differentiating that function 'f' with respect to x, treating, y and z as constants.

hartnn (hartnn):

\(\partial f/\partial y\) means you are differentiating that function 'f' with respect to y, treating, x and z as constants. \(\partial f/\partial z\) means you are differentiating that function 'f' with respect to z, treating, y and x as constants.

hartnn (hartnn):

any specific function you wanna partially differentiate ?

hartnn (hartnn):

or should i create an example ?

OpenStudy (anonymous):

Let say f(s,y,z) = 3x^3 y^2z^4

OpenStudy (anonymous):

its f(x,y,z)

OpenStudy (anonymous):

Then find fsubx (1,2,3)

hartnn (hartnn):

so for ∂f/∂x , you treat y and z as constant f = constant * x^3 where constant = 3y^2z^4

hartnn (hartnn):

can you differentiate f = constant * x^3

OpenStudy (anonymous):

Not really. I'm sorry.

hartnn (hartnn):

\(d/dx(x^n)=nx^{n-1} \\ (d/dx) x^3 =... ?\)

OpenStudy (anonymous):

aa. I know that :)

OpenStudy (anonymous):

3x^2

hartnn (hartnn):

so, \((d/dx)constant*x^3=constant *(3x^2)=3y^2z^4 *(3x^2)=...?\)

OpenStudy (anonymous):

9x^2y^2^2z^4

hartnn (hartnn):

thats your ∂f/∂x

OpenStudy (anonymous):

What about, when you have to find the slope of the surface in the x-direction at the point (2,0)?

hartnn (hartnn):

slope at a point = derivative at that point so differentiate the function once and then put x=2,y=0

OpenStudy (anonymous):

okay thanks.

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

where you're from?

hartnn (hartnn):

i am from india :) what about you ?

OpenStudy (anonymous):

Philippines.

hartnn (hartnn):

nice meeting you :)

OpenStudy (anonymous):

likewise :)

OpenStudy (anonymous):

Do you know the contour plotting in a three dimensional into two dimensional?

hartnn (hartnn):

i am afraid not, sorry :(

OpenStudy (anonymous):

Its okay :)